OpenStudy (anonymous):

Very each identiy: tan^2x/secx=secx-cosx

OpenStudy (***[isuru]***):

ok ! before u solve this one u have to remember this 1 + tan^2 x = sec^2 x then tan^2 x = sec^2 x -1 then substitute that in to the above equation.. tan^2 x / secx = (sec^x - 1)/sec x = (sec^2 x/secx ) - (1/secx) = sec x - cos x thats it! hope this will help ya!!

OpenStudy (anonymous):

Thank you so much!!!!!! :DDD